OpenStudy (anonymous):

Determine if the expression is a perfect square and if it is, factor it. 16a2 + 8a + 1 A. not a perfect square B. (a + 4)(a – 4) C. (a – 4)(a – 4) D. (4a + 1)2

OpenStudy (anonymous):

here is a hint to get you started ^_^ a perfect square must have: - two terms that are squares - and the remaining term must be 2*A*B can you pick out the two terms in your problem that are squares?

OpenStudy (anonymous):

B @DemolisionWolf

OpenStudy (anonymous):

its either B or C

OpenStudy (anonymous):

mm... not quite I'll show ya! 16a2 + 8a + 1 =(4a)^2 + 8a + 1^2 now, see how 4a is like A and 1 is like B in this formating? A^2 + 2AB + B^2 that above equals = (A+B)^2 so, lets substitute our 'A' and 'B' in if A = 4a and B = 1, then we would get (4a + 1)^2

OpenStudy (anonymous):

Factor. v2 – 36 A. (v – 6)(v – 6) B. (v + 6)(v – 6) C. (v2 – 6)2 D. (v – 6)2

OpenStudy (anonymous):

v^2 – 36 36 is a square, it is 6^2 =v^2 - 6^2 next, use the difference of squares method to solve this one: it looks like this.. A^2 - B^2 =(A - B) (A + B)

OpenStudy (anonymous):

what was the answer to the last one afterall?